the smaller of two numbers is 10 less than 4times the greater number.the greater number is 19 more than the smaller.find the two numbers.
Answer by ankor@dixie-net.com(22740)


Let x = the smaller number & y = the larger
:
the smaller of two numbers is 10 less than 4 times the greater number.
x = 4y - 10
:
the greater number is 19 more than the smaller.
y = x + 19
:
find the two numbers.
:
Substitute (4y-10) for x in the above equation
y = (4y-10) + 19
y - 4y = 19 - 10
-3y = +9
y = 9/-3
y = -3, the larger number
:
x = 4y - 10
x = 4(-3) -10
x = -12 - 10
x = -22, the smaller number
:
:
Check solution in the 2nd equation:
y = x + 19
-3 = -22 + 19